OSMOS Producing Low Image Counts

Tuesday, August 27, 2019 1:38 PM

Problem(s) Encountered:

The problem of taking long-ish exposures that turn out looking like very short 10-20 second long exposures, like that happened on the previous night, existed tonight too.
On a partially cloud night, like the last two nights, one might think its due to clouds or cirrus. But by rebooting the IE computer then restarting Prospero, the problem went away and the
resulting images look fine.
However, once things got working again, I found that any movement of the telescope, say to another object, or even sometimes doing a small shift in position, would cause the problem to reappear.  Consequently, my the end of the night, I always rebooted IE and Prospero before starting a fresh set of exposures.

Lost several hours tonight and much of the previous night to
this problem. 


It appears that the issue stemmed from a failing solenoid coil used for shutter actuation.  Since I did not have a spare on-hand, I pulled apart the Buckeye filter wheel (FW) and removed its similar solenoid.  I then pulled the solenoid from OSMOS and installed the one from the FW.  There was one issue and one bit of luck.  The issue is that the FW solenoid has two mounting holes on one side while in OSMOS, there is only one on each side.  So installing, I could only bolt it down with one bolt.  Still, that bolt is nice and tight and I do not think the solenoid is going anywhere.  Luckily, I was able to simply slide the shafts out from each solenoid, so the only part replaced is the coil.  That made things a fair amount easier.  Took a while to adjust the positioning to ensure that the shutter fully opens, but ultimately got there.  Tests at various positions appear to work as expected, so OSMOS should be back in action for tonight.Since this solenoid is different from those for the FW, I am going to search some and see if I can order a 1:1 replacement for it.  It would be nice to get it back together exactly as it was before this failure.  Also, have some breathing room now as the 2.4m FW isn’t scheduled for use at all this semester.
Telescope has been used since the repair and counts look nice and stable.